S. dx X V16-9x2

Choose the best answer:

In|csc cot | + C 2

n|csce cot | +C 4

In|csc 8 – cotoſ +C 3

In|csc – cot 0 + c 4
Answer #1

da arib-qaz 1 - 3 Let 3a=48ino #= 4 sind 3 -18 4sing da=400 COSO do 16-9213= 16-16sin?a= 1621-sina)= 16 case ~16-923 - Siocos